# Think Twice: Branch-and-Rethink Reasoning Reward Model

### v2

@@ Abstract
Large language models (LLMs) increasingly rely on thinking models that externalize intermediate steps and allocate extra test-time compute, with think-twice strategies showing that a deliberate second pass can elicit stronger reasoning. In contrast, most reward models (RMs) still compress many quality dimensions into a single scalar in one shot, a design that induces judgment diffusion: attention spreads across evaluation criteria, yielding diluted focus and shallow analysis. We introduce branch-and-rethink (BR-RM), a two-turn RM that transfers the think-twice principle to reward modeling. Turn 1 performs adaptive branching, selecting a small set of instance-critical dimensions (such as factuality and safety) and sketching concise, evidence-seeking hypotheses. Turn 2 executes branch-conditioned rethinking, a targeted reread that tests those hypotheses and scrutinizes only what matters most. We train with GRPO-style reinforcement learning over structured two-turn traces using a simple binary outcome reward with strict format checks, making the approach compatible with standard RLHF pipelines. By converting all-at-once scoring into focused, second-look reasoning, BR-RM reduces judgment diffusion and improves sensitivity to subtle yet consequential errors while remaining practical and scalable. Experimental results demonstrate that our model achieves state-of-the-art performance on three challenging reward modeling benchmarks across diverse domains. The code and the models will be released soon.
@@ 1 Introduction
Extensive experiments on reward modeling benchmarks show that our method, BR-RM, achieves overall state-of-the-art performance across RewardBench (Lambert et al., 2025b), RMBench (Liu et al., 2025b), and RMB (Zhou et al., 2025), outperforming strong baselines across multiple domains, including reasoning, general knowledge, safety, and alignment with human preferences. Beyond final performance, we conduct extensive empirical analyses of BR-RM, including the ablation study of our training recipes, the detailed analysis of the reward designs, and concrete impacts of training data sources. These insights provide a deeper understanding of reward reasoning processes and will hopefully inspire the development of future reasoning reward models.
@@ 3.2.2 Stage 2: Conditioned Rethinking for Deep Analysis
The second stage is designed to prevent shallow analysis. With the analytical scope established by $\tau_{1}$, the model performs a targeted re-evaluation. This Branch-Conditioned Rethinking is not an independent assessment but a deliberate critique of the responses through the lens of the criteria identified in Stage 1. To ensure this deliberation is principled, we incorporate task-specific evaluation hierarchies, $\mathcal{H}_{\text{task}}$, which provide a structured inductive bias. These hierarchies encode a stable ordering of importance for different criteria, such as:
  • Accuracy-Critical Tasks: $\text{Correctness}>\text{Methodology}>\text{Clarity}$
  • Creative Tasks: $\text{Intent Alignment}>\text{Quality}>\text{Novelty}$
@@ 3.3.1 The GRPO Training Algorithm
We employ Generalized Reward Policy Optimization (GRPO, Shao et al. 2024), a variant of PPO, chosen for its training stability and sample efficiency. The training loop (Algorithm 1) optimizes the policy $\pi_{\theta}$ to maximize a composite reward function over the distribution of generated traces.
@@
$L_{\text{GRPO}}=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\min\left(\rho(\tau)A(\tau),\text{clip}(\rho(\tau),1-\epsilon,1+\epsilon)A(\tau)\right)\right]-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})$ (3)
@@ 4.1 Experimental Setup
Implementation Details. Our implementation leverages a scalable and efficient post-training library, NeMo-RL [^1] https://github.com/NVIDIA-NeMo/RL, with a custom two-turn reward modeling environment. Regarding model design, the penalty constant for format checking is -100 and the weight for reward terms is 10. For training, model-specific learning rates were tuned based on scale: $5\times 10^{-7}$ for Qwen3-8B models (Team, 2025) and $1\times 10^{-6}$ for Qwen3-14B (Team, 2025) models, all trained for 400 optimization steps with validation every 10 steps. For inference, we employ a consistent evaluation protocol with training parameters, computing accuracy as the primary metric, with additional domain-specific metrics for comprehensive analysis. More configurations are in Appendix D.
Table 1: The performance comparison between best-performing baselines. Bold numbers indicate the best performance, Underlined numbers indicate the second best. The more detailed numbers on RewardBench, RM-Bench, and RMB are in Table 6, Table 7, and Table 8 in the Appendix.
@@
SteerLM-RM-70B 88.8 52.5 58.2 66.5
Eurus-RM-7b 82.8 65.9 68.3 72.3
Internlm2-20b-reward 90.2 68.3 62.9 73.6
Skywork-Reward-Gemma-2-27B 93.8 67.3 60.2 73.8
Internlm2-7b-reward 87.6 67.1 67.1 73.9
ArmoRM-Llama3-8B-v0.1 90.4 67.7 64.6 74.2
Nemotron-4-340B-Reward 92.0 69.5 69.9 77.1
Skywork-Reward-Llama-3.1-8B 92.5 70.1 69.3 77.5
INF-ORM-Llama3.1-70B 95.1 70.9 70.5 78.8
@@
Claude-3.5-sonnet-20240620 84.2 61.0 70.6 71.9
Llama3.1-70B-Instruct 84.0 65.5 68.9 72.8
Gemini-1.5-pro 88.2 75.2 56.5 73.3
Skywork-Critic-Llama-3.1-70B 93.3 71.9 65.5 76.9
GPT-4o-0806 86.7 72.5 73.8 77.7
@@
JudgeLRM-7B 75.2 64.7 53.1 64.3
DeepSeek-GRM-27B 86.0 72.7 69.0 75.9
Self-taught-evaluator-Llama3.1-70B 90.2 71.4 67.0 76.2
RM-R1-Qwen-Instruct-14B 88.2 76.1 69.2 77.8
RM-R1-DeepSeek-Distilled-Qwen-14B 88.9 81.5 68.5 79.6
EvalPlanner-Llama-3.3-70B-Instruct 93.8 82.1 - -
J1-Llama-70B 93.3 82.7 - -
RM-R1-Qwen-Instruct-32B 91.4 79.1 73.0 81.2
RM-R1-DeepSeek-Distilled-Qwen-32B 90.9 83.9 69.8 81.5
RRM-32B 91.2 84.0 70.0 81.7
@@ 4.2 Main Results
Table 1 presents a comprehensive performance comparison of our Branch-and-Rethink Reward Models (BR-RMs) against the best-performing baselines across three standard benchmarks.
ScalarRMs remain competitive but plateau. ScalarRMs such as INF-ORM-Llama3.1-70B achieve the strongest RewardBench score (95.1) and lead most scalar baselines in average accuracy (78.8). However, despite high scores on RewardBench, they lag behind reasoning-focused models on RM-Bench (70.9) and RMB (70.5). Smaller ScalarRMs perform relatively well, suggesting that simply scaling up parameters without reasoning mechanisms brings diminishing returns.
@@ 4.3 Ablation Study
Prompting strategy BR-RM-Qwen-8B BR-RM-Qwen-14B
RB RMBench RMB Avg RB RMBench RMB Avg
BR-RM (Full Model) 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
Branching Only 90.1 82.6 67.4 80.0 90.9 83.2 71.0 81.7
Unconditioned Rethink 90.6 83.9 69.3 81.3 91.3 84.4 71.9 82.5
Single-Turn Model 90.4 84.5 70.0 81.6 91.3 85.2 73.3 83.3
To validate the contributions of our proposed components, we conduct several ablations based on our BR-RM models. We systematically remove key architectural components to isolate their impact on performance across our evaluation benchmarks (Table 2).
Branching Only In this variant, we remove the second “rethinking” turn. This ablation caused the most significant performance degradation. The first turn is designed for both branching and error detection, which directs its focus toward identifying specific, granular issues. Without the second turn, the model’s final judgment relies entirely on this initial pass. This can mislead the model to over-emphasize details and minor flaws without having a comprehensive view of the whole response. The absence of the holistic re-evaluation step is particularly damaging on nuanced benchmarks like RMB, where the 14B model’s performance drops by a substantial 3.7 points. This shows that the second turn is critical for contextualizing initial findings.
Unconditioned Rethinking Here, we retain the two-turn training but remove branching. The first turn goes through all the dimensions while the second turn is prompted with a generic “re-evaluate the response” instruction. While this model performs better than the “Branching Only” version, it remained well below our full model. For instance, the 14B model sees performance drops of 0.8, 1.5, and 2.8 points on the respective benchmarks. This result underscores the importance of adaptive focus. Without an explicit directive from the first turn, the second pass is less efficient and prone to the same judgment diffusion it is designed to prevent.
Single-Turn Model To test the value of our sequential process, we merge the prompts for both turns into a single instruction. This approach forces the model to rely on only one-time thinking. The resulting prompt is highly complex, demanding that the model handles too many subtasks at the same time: select dimensions, sketch issues, and then re-evaluate to produce a final score. This cognitive overload prevents the model from deeply engaging with any single subtask, leading to a shallower analysis. The consistent, albeit smaller, performance drop confirms that a deliberate, sequential process is critical for high-quality, reasoned evaluation.
@@ 4.4 Reward Design Analysis
Reward BR-RM-Qwen-8B BR-RM-Qwen-14B
RB RMBench RMB Avg RB RMBench RMB Avg
Our Reward 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
Removing Format Checks 90.5 84.2 71.5 82.1 91.3 85.4 73.1 83.3
Scoring on a Scale 90.0 83.5 70.1 81.2 90.8 85.1 71.7 82.5
Additional Reward for Branch 86.5 79.0 66.2 77.2 87.8 82.7 69.3 79.9
Scoring on a Scale In another experiment, we had the model output a detailed score on a scale, like from -3 to 3. This is inspired by the reward dataset, HelpSteer3, whose human-annotated preferences use a similar scale to express preference strength. During training, the reward was based on the distance between the predicted score and the ground-truth score. To process other datasets with only binary labels, we arbitrarily mapped the winning response to +2 and the losing one to -2. This approach failed, possibly due to a disconnect between the training task and the final evaluation: while training rewarded the model for predicting an exact score, our ultimate evaluation only cared about whether the model correctly identified the winning response (a simple binary choice). This means that the model had no strong incentive to learn the true meaning behind the scores; it only needed to learn if a score was positive or negative. Predicting precise scores is less beneficial when only the winner matters.
@@
The Best Reward: Binary Reward with Format Checking. After testing these alternatives, we confirmed that the best reward design is the most straightforward one. In this setup, the model produces a binary choice (A or B) within a required format and receives a simple reward based on whether it is correct. Each of the other designs introduced a flaw – confusing signals, misaligned objectives, or indirect feedback – that hurt performance. This shows that a direct alignment between the task and a reward based on the correct answer provides the most effective foundation for training.
4.5 Training Data Analysis
Model BR-RM-Qwen-8B BR-RM-Qwen-14B
RB RMBench RMB Avg RB RMBench RMB Avg
Full training data 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
w/o HelpSteer3 90.4 83.5 68.0 80.6 91.5 84.4 70.3 82.1
w/o Skywork 90.6 84.2 66.1 80.3 91.3 85.1 69.7 82.0
w/o Math-Step-DPO 91.0 84.7 70.7 82.1 91.5 85.0 72.6 83.0
w/o Code-Preference 91.1 84.8 71.2 82.4 91.8 85.4 74.0 83.7
To understand each dataset’s contribution, we conducted an ablation study, systematically removing one at a time to measure its impact on model performance. From Table 4, this analysis revealed a clear distinction between the roles of foundational and specialized datasets. For HelpSteer3, removing this foundational dataset led to a broad performance decline, with a 4.4-point drop on the RMB benchmark and a 1.5-point drop on RMBench. This highlights its critical role in establishing general understanding and factual accuracy. For Skywork, excluding our only dataset with safety data caused the most severe impact—a 5.0-point drop on the safety-focused RMB benchmark. This confirms that explicit safety training is essential for model harmlessness. For Math-Step-DPO, removing this specialized dataset predictably weakened reasoning, causing a 2.1-point drop on RMB and a 0.9-point drop on RMBench. This shows in-domain data is necessary for complex, multi-step reasoning. For Code-Preference, the exclusion of this smaller dataset resulted in the most modest decrease (0.7 points on RMB and 0.5 points on RMBench). This suggests its logical reasoning contributions are partially covered by the broader datasets.

### v3

@@ Abstract
Large language models (LLMs) increasingly rely on thinking models that externalize intermediate steps and allocate extra test-time compute, with think-twice strategies showing that a deliberate second pass can elicit stronger reasoning. In contrast, most reward models (RMs) still compress many quality dimensions into a single scalar in one shot, a design that induces judgment diffusion: attention spreads across evaluation criteria, yielding diluted focus and shallow analysis. We introduce branch-and-rethink (BR-RM), a two-turn RM that transfers the think-twice principle to reward modeling. Turn 1 performs adaptive branching, selecting a small set of instance-critical dimensions (such as factuality and safety) and sketching concise, evidence-seeking hypotheses. Turn 2 executes branch-conditioned rethinking, a targeted reread that tests those hypotheses and scrutinizes only what matters most. We train with GRPO-style reinforcement learning over structured two-turn traces using a simple binary outcome reward with strict format checks, making the approach compatible with standard RLHF pipelines. By converting all-at-once scoring into focused, second-look reasoning, BR-RM reduces judgment diffusion and improves sensitivity to subtle yet consequential errors while remaining practical and scalable. Experimental results demonstrate that our model achieves state-of-the-art performance on three challenging reward modeling benchmarks across diverse domains.
@@ 1 Introduction
Extensive experiments on reward modeling benchmarks show that our method, BR-RM, achieves overall state-of-the-art performance across RewardBench (Lambert et al., 2025b), RMBench (Liu et al., 2025b), and RMB (Zhou et al., 2025), outperforming strong baselines across multiple domains, including reasoning, general knowledge, safety, and alignment with human preferences. Beyond final performance, we conduct extensive empirical analyses, including the ablation study of our training recipes, the detailed analysis of the reward designs, and concrete impacts of training data sources. These insights provide a deeper understanding of reward reasoning processes and will hopefully inspire the development of future reasoning reward models.
@@ 3.2.2 Stage 2: Conditioned Rethinking for Deep Analysis
The second stage is designed to prevent shallow analysis. With the analytical scope established by $\tau_{1}$, the model performs a targeted re-evaluation. This Branch-Conditioned Rethinking is not an independent assessment but a deliberate critique of the responses through the lens of the criteria identified in Stage 1. To ensure this deliberation is principled, we incorporate task-specific evaluation hierarchies $\mathcal{H}_{\text{task}}$ to provide a structured inductive bias. For example, accuracy-driven tasks prioritize Correctness and Methodology over Clarity, whereas creative tasks favor Intent Alignment and Novelty.
@@ 3.3.1 The GRPO Training Algorithm
We employ Generalized Reward Policy Optimization (GRPO, Shao et al. 2024), a variant of PPO, chosen for its training stability and efficiency. The training loop (Algorithm 1) optimizes the policy $\pi_{\theta}$ to maximize a composite reward function over the distribution of generated traces.
@@
$L_{\text{GRPO}}={}\mathbb{E}_{\tau\sim\pi_{\theta}}\Big[\min\Big(\rho(\tau)A(\tau),$ (3)
$\quad\text{clip}(\rho(\tau),1-\epsilon,1+\epsilon)A(\tau)\Big)\Big]-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})$
@@ 4.1 Experimental Setup
Implementation Details. Our implementation leverages a scalable and efficient post-training library, NeMo-RL [^1] https://github.com/NVIDIA-NeMo/RL, with a custom two-turn reward modeling environment. Regarding model design, the penalty constant for format checking is -100 and the weight for reward terms is 10. For training, model-specific learning rates were tuned based on scale: $5\times 10^{-7}$ for Qwen3-8B models (Team, 2025) and $1\times 10^{-6}$ for Qwen3-14B (Team, 2025) models, all trained for 400 optimization steps. For inference, we employ a consistent evaluation protocol with training parameters, computing accuracy as the primary metric, with additional domain-specific metrics for comprehensive analysis. More configurations are in Appendix D.
Table 1: The performance comparison between best-performing baselines. Bold and underline indicate the first and second best performance respectively.
@@
SteerLM-RM-70B (Chen et al., 2025b) 88.8 52.5 58.2 66.5
Eurus-RM-7b (Yuan et al., 2025) 82.8 65.9 68.3 72.3
Internlm2-20b-reward (Cai et al., 2024) 90.2 68.3 62.9 73.6
Skywork-Reward-Gemma-2-27B (Liu et al., 2024b) 93.8 67.3 60.2 73.8
Internlm2-7b-reward (Cai et al., 2024) 87.6 67.1 67.1 73.9
ArmoRM-Llama3-8B-v0.1 (Chen et al., 2025b) 90.4 67.7 64.6 74.2
Nemotron-4-340B-Reward (Adler et al., 2024) 92.0 69.5 69.9 77.1
Skywork-Reward-Llama-3.1-8B (Liu et al., 2024b) 92.5 70.1 69.3 77.5
INF-ORM-Llama3.1-70B (Chen et al., 2025b) 95.1 70.9 70.5 78.8
@@
Claude-3.5-sonnet (Anthropic, 2024) 84.2 61.0 70.6 71.9
Llama3.1-70B-Instruct (Dubey et al., 2024) 84.0 65.5 68.9 72.8
Gemini-1.5-pro (Reid et al., 2024) 88.2 75.2 56.5 73.3
Skywork-Critic-Llama-3.1-70B (Liu et al., 2024b) 93.3 71.9 65.5 76.9
GPT-4o-0806 (OpenAI, 2023) 86.7 72.5 73.8 77.7
@@
JudgeLRM-7B (Chen et al., 2025a) 75.2 64.7 53.1 64.3
DeepSeek-GRM-27B (Liu et al., 2025c) 86.0 72.7 69.0 75.9
Self-taught-evaluator-70B (Wang et al., 2024) 90.2 71.4 67.0 76.2
RM-R1-Qwen-14B (Chen et al., 2025b) 88.2 76.1 69.2 77.8
RM-R1-DS-Distill-14B (Chen et al., 2025b) 88.9 81.5 68.5 79.6
EvalPlanner-Llama-3.3-70B (Saha et al., 2025) 93.8 82.1
J1-Llama-70B (Whitehouse et al., 2025) 93.3 82.7
RM-R1-Qwen-32B (Chen et al., 2025b) 91.4 79.1 73.0 81.2
RM-R1-DS-Distill-32B (Chen et al., 2025b) 90.9 83.9 69.8 81.5
RRM-32B (Guo et al., 2025) 91.2 84.0 70.0 81.7
@@ 4.2 Main Results
Table 1 compares BR-RMs against state-of-the-art baselines across three benchmarks.
ScalarRMs plateau on reasoning tasks. While models like INF-ORM-Llama3.1-70B dominate RewardBench (95.1), they lag significantly on reasoning-heavy benchmarks (RM-Bench 70.9, RMB 70.5) , suggesting diminishing returns from parameter scaling alone.
GenRMs trade off consistency. Generative models improve robustness but fail to consistently surpass scalars. Skywork-Critic excels on RewardBench yet struggles on RMB (65.6), and GPT-4o, despite a balanced profile, still trails leading scalar and reasoning RMs.
ReasonRMs require scale. Reasoning mechanisms show clear gains only beyond 14B parameters. The 32B RM-R1 models achieve strong results (up to 83.9 on RM-Bench) , whereas smaller reasoning models (e.g., 7B) underperform even weaker ScalarRMs, indicating that reasoning requires sufficient model capacity.
BR-RM achieves the best overall balance. Our BR-RM-Qwen-14B attains the highest average score (84.2), setting a new SOTA on RM-Bench (85.9) and ranking top-2 on RMB (74.7). Notably, our 8B model outperforms significantly larger baselines (e.g., GPT-4o, 70B scalars), confirming that adaptive two-turn reasoning sets a new bar for both efficiency and robustness.
@@ 4.3 Ablation Study
Figure 3: Ablation Study. The performance difference from the full models highlighting the contribution of each component. RewardBench is abbreviated as RB.
To validate the contributions of our proposed components, we conduct several ablations based on our BR-RM models. We systematically remove key architectural components to isolate their impact on performance across our evaluation benchmarks (Figure 3).
Branching Only Removing the second "rethinking" turn caused the most significant degradation, with a 3.7-point drop on RMB for the 14B model. Without the second turn, the model over-emphasizes granular flaws detected in the first pass without a comprehensive perspective. This confirms that the holistic re-evaluation step is critical for contextualizing initial findings.
Unconditioned Rethinking We tested a version with two turns but no branching, using a generic re-evaluation prompt for the second pass. Performance dropped by up to 2.8 points, demonstrating that without an explicit directive from the first turn, the second pass suffers from judgment diffusion and reduced efficiency. Adaptive focus is essential for localized error correction.
Single-Turn Model Merging both turns into a single instruction forces the model to perform dimension selection, error sketching, and final scoring simultaneously. This cognitive overload leads to shallower analysis compared to our sequential process. The resulting performance decline confirms that a deliberate, multi-step pipeline is necessary for high-quality evaluation.
@@ 4.4 Reward Design Analysis
Figure 4: Reward Analysis. Quantitative comparison of our proposed reward design against several ablations. Our approach using a direct binary reward consistently achieves the highest scores. Here RewardBench is abbreviated as RB.
To identify the best training strategy, we evaluated several reward designs in Figure 4.
Removing Format Checks Removing strict formatting allowed the model to find shortcuts, skipping reasoning to produce only the final answer. Format checks provide essential structure and help the model distinguish between incorrect reasoning and poor formatting. Without this control, ambiguous feedback slows down learning. Format checks ensure structural integrity and clear, unambiguous feedback.
Scoring on a Scale We tested a scoring scale to reflect preference strength, mapping binary labels to the range from -3 to 3. This failed because the model focused on predicting exact values rather than the relative quality needed to identify a winner. This misalignment between the training task and the final evaluation objective hindered performance. Predicting precise scores is less beneficial when only the win/loss outcome matters.
Additional Reward for Branching We attempted to reward the first reasoning step using indirect values calculated from the branching process. However, this lacked a natural ground-truth signal and created two problems: computational costs nearly doubled, and the resulting reward signals were too infrequent and erratic to stabilize training. Calculated intermediate rewards generate noisy and unstable training signals.
@@
The Best Reward: Binary Reward with Format Checking The most straightforward design—binary choice within a required format—proved most effective. Other designs introduced confusing signals or misaligned objectives that hurt performance. Direct alignment between the task and a reward based on the correct answer provides the most effective training foundation.
4.5 Reasoning Turn Scaling
Figure 5: Ablation on Turn Scaling (Qwen-14B). The trend lines across three benchmarks (RewardBench, RM-Bench, RMB) illustrate that performance consistently peaks at the 2-turn mark and declines with further rethinking, highlighting the diminishing returns of excessive reasoning turns.
We explore the optimality of the two-turn design by scaling the number of reasoning turns. We trained variants of our 14b model with 1, 3, and 4 turns, where additional turns involve iterative rethinking.
As shown in Figure 5, performance peaks at 2 turns and degrades with additional iterations. We hypothesize that beyond two turns, the model begins to “over-think,” hallucinating subtle errors during repeated verification or drifting from the original context. This confirms that a single hypothesis generation step followed by a verification step is the most robust configuration, striking an optimal balance between reasoning depth and stability.
4.6 Distribution of Selected Branches
Figure 6: Distribution of Selected Branches across RM-Bench Subsets. The heatmap visualizes the sparse and domain-specific activation patterns, demonstrating that the model adaptively prioritizes relevant dimensions (e.g., Logical Reasoning for Math, Safety Awareness for Safety) rather than applying a uniform rubric.
Table 2: Inference Efficiency. Latency is measured per instance on $2\times$ H100 GPUs. Our two-turn 8B model outperforms the larger 32B single-turn baseline in both speed and accuracy.
Model Type Latency (s) Avg Acc (%)
InternLM2-7B-Reward Scalar 0.015 72.3
INF-ORM-Llama3.1-70B Scalar 0.25 78.8
RM-R1-Qwen-Instruct-7B Gen (1-Turn) 5.1 73.9
RM-R1-Qwen-Instruct-32B Gen (1-Turn) 10.3 81.2
BR-RM-Qwen-8B (Ours) Gen (2-Turn) 9.5 82.6
BR-RM-Qwen-14B (Ours) Gen (2-Turn) 15.9 84.2
To investigate whether predefined dimensions limit generality, we analyzed branch selection behavior on RM-Bench, observing sparse and adaptive activation (averaging 2.5 dimensions per instance) where specific dimensions like Safety Awareness” or Implementation Capability” heavily dominate their respective domains (91% and 94% activation, respectively). This confirms that our universal set functions as a flexible semantic basis rather than a rigid checklist, allowing the model to dynamically map novel failure modes into high-level cognitive categories and effectively prune the search space to prevent focus dilution.
4.7 Inference Efficiency
A primary concern with reasoning-based reward models is the increased inference cost compared to traditional scalar models. To quantify this trade-off, we conducted a controlled latency benchmark on $2\times$ H100 GPUs. While scalar RMs are inherently faster (e.g., InternLM2-7B-Reward takes 0.015s), they hit a distinct performance ceiling, peaking at 78.8% average accuracy with INF-ORM-Llama3.1-70B. In the regime of reasoning-based evaluation, our Branch-and-Rethink strategy proves significantly more efficient than simply scaling model parameters. As shown in Table 2, ourBR-RM-Qwen-8B requires 9.5s per instance, making it faster than the single-turn baseline RM-R1-Qwen-Instruct-32B (10.3s) while achieving higher average accuracy (82.6% vs. 81.2%). It confirms that allocating test-time compute to a targeted second pass yields better than generating a single broad critique with a larger model.
Beyond raw latency, the practical feasibility of BR-RM should be viewed through the lens of high-stakes alignment pipelines. In offline reinforcement learning settings like Rejection Sampling or PPO, the quality of the reward signal is often the primary bottleneck rather than generation speed; a noisy reward signal from a faster but less accurate scalar model can destabilize training. Furthermore, the inference latency of BR-RM is effectively amortized in modern asynchronous RL setups, where reward generation is decoupled from policy updates. This architecture prevents the reward model from becoming a blocking factor in the training loop, justifying the investment in “thinking twice” for superior preference labeling.